.          A 50.0-mL sample of gas is cooled from 119°C to 80.0°C. If the pressure remains constant, what is the final volume of the gas?


1
Expert's answer
2021-03-09T06:15:24-0500

V1 = 50.0 mL

T1 = 119°C + 273 = 392 K

P1 = P2 = const

V2 = unknown

T2 = 80.0°C + 273 = 353 K


Solution:

Since the pressure doesn't change, Charles' Law can be used.

Charles' Law can be expressed as: V1 / T1 = V2 / T2

V1T2 = V2T1

To find the final volume, solve the equation for V2:

V2 = V1T/ T1

V2 = (50.0 mL × 353 K) / (392 K) = 45.0255 mL = 45.0 mL

V2 = 45.0 mL


Answer: The final volume of the gas is 45.0 mL.